OpenStudy (anonymous):

PLEASE HELP WITH TWO ALGEBRA PROBLEMS FOR MEDALS!!!! QUESTION 1: In a classroom, 11 out of 17 students surveyed say they had two or more siblings. If the entire school has 850 students, predict the number of students that have two or more siblings. ___ students QUESTION 2: The ratio of the perimeters of two rectangles is 4 to 9. If the perimeter of the larger rectangle is 54 inches, what is the perimeter of the smaller rectangle? ____ inches

OpenStudy (aripotta):

well the first one is almost exactly like the other one i helped you with. maybe you could do that one on your own?

OpenStudy (anonymous):

\[\frac{ 11 }{ 17 }=\frac{ x }{ 850 }\] then, 17x = 11 x 850 17x = 9350 9350 divded by 17 is 550

OpenStudy (aripotta):

yep

OpenStudy (anonymous):

ok i think i got the next one :) do you want a medal?

OpenStudy (aripotta):

yess :3 you wanna post your answer and i'll tell you if you got it right?

OpenStudy (anonymous):

Okay here it is, \[\frac{ 4 }{ 9 }=\frac{ x }{ 54 }\] then I used cross multiplication, 9x = 4 x 54 9x = 216 216 dvided by 9 is 24 is that right?

OpenStudy (aripotta):

yes! \(\Large \color{magenta}{\ddot\smile} \)

OpenStudy (anonymous):

Awesome! Thanks so much for your help @AriPotta!

OpenStudy (aripotta):

\(\Large \bf \color{#008AFF}{n}\color{#0073FF}{o} ~\color{#005CFF}{p} \color{#0045FF}{r} \color{#0017FF}{o} \color{#0200FF}{b} \color{#3000FF}{l} \color{#4700FF}{e} \color{#7500FF}{m} \color{#8C00FF}{o}~\LARGE\color{#D100FF}{\ddot\smile}\)